tan2A =3/4 then find value of tanA and express it on other ratios

Grade:10

1 Answers

Arun
5 years ago
As we know
Tan2A = 2 tan A/1 - tan² A
Now
3/4 = 2 tan A/1 - tan²A
3 - 3 tan²A = 8 tan A
3 tan² A + 8 tanA -3 = 0
(tan A+ 3) (3 tanA -1) = 0
Now you have find tanA hence you can find other ratios too